OpenStudy (anonymous):

How do you solve for tanxsin^2x=tanx?

OpenStudy (anonymous):

x between 0 and 2pi

OpenStudy (anonymous):

\[\tan x ARCsin x=\]

OpenStudy (anonymous):

Since you have tan x on both sides, you can divide each side by tan x to simplify to sin^2 x = 1. Then square each side: sin x = 1, assuming that you don't need the negative square root. Solve sin x = 1: what angle measure gives you that?

OpenStudy (anonymous):

Not arc sin, sorry sin^2x

OpenStudy (anonymous):

Ok, so you can just get rid of the tan?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

and no, you can't get rid of the tan you will lose the obvious solution \(\tan(x)=0\) if you do

OpenStudy (anonymous):

Wait you're right! sorry lol

OpenStudy (anonymous):

So one answer is 90 but I have to find another?

OpenStudy (anonymous):

\[\tan(x)\sin^2(x)=\tan(x)\]\[\tan(x)\sin^2(x)-\tan(x)=0\] \[\tan(x)\left(\sin^2(x)-1\right)=0\] etc

OpenStudy (anonymous):

Then you can change sin^2x-1 to cos^2x? But then what do you do with tanxcosx or is that wrong?

OpenStudy (anonymous):

Ohh, do you set each equal to 0?